OpenStudy (anonymous):

Find the limit of the function algebraically. limit as x approaches nine of quantity x squared minus eighty one divided by quantity x minus nine.

OpenStudy (campbell_st):

so the problem looks like \[\lim_{x \rightarrow 9} \frac{x^2 - 81}{x - 9}\] is that correct..?

OpenStudy (anonymous):

\[\lim_{x \rightarrow 9} \frac{ x ^{2}-81 }{ x-9 }\]

OpenStudy (solomonzelman):

use: \(\large\color{black}{ a^2-b^2=(a-b)(a+b) }\) for the top

OpenStudy (campbell_st):

so factor the numerator, its the difference of 2 squares.... then you'll see a common factor to remove. and finally substitute x = 9 into whats left to find the limit

OpenStudy (anonymous):

yes, well is the answer 18? I did (x+9)(x-9)/(x-9), simplified it to (x+9) and substituted 9 for x and added and got 18...

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

that is right:)

OpenStudy (anonymous):

Ok thanks so much:)) you rock:))

OpenStudy (campbell_st):

well done

OpenStudy (solomonzelman):

you welcome!